Find the integral curves of the following equations:dx/(x^2–y^2–yz) = dy/(x^2–y^2–zx) = dz/[z(x–y)]
1
Expert's answer
2020-02-11T12:06:59-0500


Solution.

1) Subtract the first expression from the second expression and equate to the third

"\frac{dx-dy}{x^2-y^2-yz-(x^2-y^2-zx)}=\frac{dz}{z(x-y)}\\
\frac{d(x-y)}{x^2-y^2-yz-x^2+y^2+zx}=\frac{dz}{z(x-y)}\\
\frac{d(x-y)}{z(x-y)}=\frac{dz}{z(x-y)}\\
d(x-y)=dz\\
\int d(x-y)=\int dz\\
z=x-y+C_1."

Then the first integral curves is

"z-x+y=C_1."

2) Add the first expression to the second and equate to the third

"\frac{dx+dy}{x^2-y^2-yz+x^2-y^2-xz}=\frac{dz}{z(x-y)}\\
\frac{d(x+y)}{2(x^2-y^2)-z(x+y)}=\frac{dz}{z(x-y)}\\
\frac{d(x+y)}{2(x -y )(x+y)-z(x+y)}=\frac{dz}{z(x-y)}\\"

Because

 "x-y=z-C_1."

Then

"\frac{d(x+y)}{2(z-c_1 )(x+y)-z(x+y)}=\frac{dz}{z(z-c_1)}\\
\frac{d(x+y)}{(x+y)(2(z-c_1)-z)}=\frac{dz}{z(z-c_1)}\\
\frac{d(x+y)}{(x+y)(z-2c_1)}=\frac{dz}{z(z-c_1)}\\
\frac{d(x+y)}{(x+y)}=\frac{(z-2c_1)dz}{z(z-c_1)}\\
\frac{d(x+y)}{(x+y)}=\frac{2z-2c_1-z}{z(z-c_1)}dz\\
\frac{d(x+y)}{(x+y)}=\left(\frac{2}{z}-\frac{1}{z-c_1}\right)dz\\
\int\frac{d(x+y)}{(x+y)}=\int\left(\frac{2}{z}-\frac{1}{z-c_1}\right)dz\\
\ln|x+y|=2\ln|z|-\ln|z-c_1|+\ln|c_2|\\
x+y=\frac{z^2 c_2}{z-c_1}"

Substitute the value "c_1"

"x+y=\frac{z^2 c_2}{z-(z-x+y)}\\
x+y=\frac{z^2 c_2}{x-y}\\
x^2-y^2=z^2 c_2"

Then the second integral curves is

"\frac{x^2-y^2}{z^2}=c_2"